OpenStudy (samhang):

how to find dy/dx if \[x ^{2}y=x ^{2}+y ^{2}\]

OpenStudy (jiteshmeghwal9):

\[y^2=x^2y-x^2\]differentiate both sides with the help of chain rule

OpenStudy (jiteshmeghwal9):

\[\frac{d(y^2)}{dy}.\frac{dy}{dx}=x^2\frac{dy}{dx}+y \frac{d(x^2)}{dx}- \frac{d(x^2)}{dx}\]

OpenStudy (samhang):

is it \[\frac{ dy }{ dx}=( x ^{2}+2xy-2x)/2y\] answer

OpenStudy (jango_in_dtown):

no

OpenStudy (mathmale):

It'd be much easier to give you worthwhile feedback if you'd share ALL of your work.

OpenStudy (samhang):

how about this upload....is it Correct!

OpenStudy (jango_in_dtown):

its correct @samhang

OpenStudy (samhang):

Thanks guys